What is the solution set of x+y=7 and x-y=1?
Answer by checkley77(12569)

x+y=7
x-y=1 add.
------
2x=8
x=8/2
x=4 ans.
4+y=7
y=7-4
y=3 ans.
proof:
4-3=1
1=1